# Adaptive Text Watermark for Large Language Models

### v1

@@ Abstract
The advancement of Large Language Models (LLMs) has led to increasing concerns about the misuse of AI-generated text, and watermarking for LLM-generated text has emerged as a potential solution. However, it is challenging to generate high-quality watermarked text while maintaining strong security, robustness, and the ability to detect watermarks without prior knowledge of the prompt or model. This paper proposes an adaptive watermarking strategy to address this problem. To improve the text quality and maintain robustness, we adaptively add watermarking to token distributions with high entropy measured using an auxiliary model and keep the low entropy token distributions untouched. For the sake of security and to further minimize the watermark’s impact on text quality, instead of using a fixed green/red list generated from a random secret key, which can be vulnerable to decryption and forgery, we adaptively scale up the output logits in proportion based on the semantic embedding of previously generated text using a well designed semantic mapping model. Our experiments involving various LLMs demonstrate that our approach achieves comparable robustness performance to existing watermark methods. Additionally, the text generated by our method has perplexity comparable to that of un-watermarked LLMs while maintaining security even under various attacks. Our code will be available at github.com/yepengliu/adaptive-watermark.
@@ 1 Introduction
The rapid development of Large Language Models (LLMs) (Zhang et al., 2022; Touvron et al., 2023) has ushered in transformative possibilities but also brings forth potential challenges. One primary challenge lies in the misuse of AI-generated text for malicious purposes, such as spreading disinformation, generating fake news, or engaging in plagiarism. The sheer volume and quality of AI-generated text make it increasingly difficult to discern between human and AI-generated content. As a solution, implementing a watermark on LLMs proves to be an effective method for detecting AI-generated text.
@@
Specifically, the watermarking technique encodes certain patterns into the text generated by LLMs, which are imperceptible to humans but can be detected by the algorithm. This is usually achieved by introducing well-designed perturbations to the original logits or using a special sampling method. Kirchenbauer et al. (2023a) design a watermark for LLM by randomly separating the vocabulary into ‘green’ and ‘red’ lists and increasing the frequency of tokens from the ‘green’ list. The previously generated token is used as a random key to determine the ‘green/red’ list for the next token. This method achieves good detection performance when the watermarked text remains unmodified. However, after modifications such as word replacement and paraphrasing, this watermarking method will fail as the ‘green/red’ list cannot be reliably recovered. As it is common for individuals to alter AI-generated text rather than use it directly, improving the robustness of the watermark becomes an essential requirement.
@@
To improve the robustness of LLM watermarking, Zhao et al. (2023) propose to use a fixed ‘green/red’ list at each time step, which is generated from a random secret key that is shared with the detector. Both empirical and theoretical evidence demonstrate that this method substantially enhances robustness. More recently, Sadasivan et al. (2023) introduces a spoofing attack specifically designed to deduce hidden watermark signatures by analyzing the token frequency, thereby enabling the forgery of watermarked text. Such an attack offers a method for adversaries to circumvent the watermark, potentially damaging the reputation of the system. Regrettably, it has been shown that the watermarking algorithm proposed in Zhao et al. (2023) is vulnerable to spoofing attack, and one can infer the fixed ‘green’ token list and forge the watermarked text.
@@
The studies above enhance the frequency of tokens in the ‘green’ list during generation by uniformly adding a fixed value to the logits of those ‘green’ tokens. This way of perturbing the distribution may change the distribution significantly, deteriorating the quality of the watermarked text. Therefore, Kuditipudi et al. (2023) presents a distortion-free watermarking method that maintains the original probability distribution of LLMs and embeds the watermark during the token sampling process. This method provides a watermarked text whose text perplexity is close to that of un-watermarked text. However, such a distortion-free watermarking lacks sufficient robustness against modifications, like paraphrasing.
@@
Inspired by the aforementioned research, designing an effective watermark algorithm requires us to comprehensively consider at least the following three factors. (1) Robustness: The watermark should remain detectable after modifications without significantly changing the semantics of the watermarked text. (2) Security: The watermark should be resilient against efforts to reverse-engineer or decryption without authorization. (3) Text quality: LLMs-generated watermarked text should preserve a level of perplexity comparable to the un-watermarked text. Among existing works, achieving high-quality text often comes with the trade-off of robustness, and prioritizing strong robustness can potentially compromise security. Therefore, designing a watermark that simultaneously guarantees robustness, security, and text quality poses a significant challenge.
@@
Figure 1: Workflow of the proposed Adaptive Watermark. Our Adaptive Watermark method will assess the entropy of the previously generated text at every time step and add watermarks only to high-entropy tokens. This figure illustrates two cases in the text generation process. For the first example, the next token distribution is high entropy as measured by AWTI, indicating high uncertainty. Then, the SLSVE module will extract the logits scaling vector based on the semantics of this text. Subsequently, we apply AWTS to perturb the sampling distribution. When it comes to the second text, it results in low entropy as measured by AWTI, suggesting a low uncertainty for the next token. In this situation, the next token will be directly sampled from the original probability distribution.
@@
In this work, our goal is to provide a holistic approach that integrates robustness and security with the production of high-quality text. Additionally, we also ensure the watermark can be detected in an agnostic manner, meaning that the watermark detection process is independent of the original LLMs and prompts. Our watermark strategy consists of three steps as shown in Figure 1.
  • [leftmargin=*,topsep=0.1em,itemsep=-0.2em]
@@
  • To minimize the impact of the watermark on the text quality while maintaining the robustness, we propose the Adaptive Watermark Token Identification (AWTI) so that the watermarking is added adaptively to high-entropy token distributions, and leaves the low-entropy token distributions intact. This method employs an auxiliary Measurement model to assess the entropy of each token, ensuring that the entropy can be reliably estimated during detection.
@@
  • To further enhance the security of our method, we introduce the Semantic-based Logits Scaling Vector Extraction (SLSVE) to replace the widely used ‘green/red’ list. Our logits scaling vector is obtained by passing the semantics of the generated text (using a pre-trained Sentence embedding model) to a Semantic Mapping Model, which is hard to reverse-engineer.
@@
We conduct extensive experiments over various language models (OPT-2.7B, OPT-6.7B (Zhang et al., 2022), GPT-J-6B (Wang & Komatsuzaki, 2021)) to show the effectiveness of our method. We employ perplexity as a metric to evaluate text quality, and our findings indicate that the quality of the watermarked text remains comparable to that of the un-watermarked text. To show the robustness of our method, we paraphrase the watermarked text using GPT3.5 and DIPPER (Krishna et al., 2023). The outcomes are on par with those achieved by Zhao et al. (2023) which is the most robust method in the literature. The spoofing attack is used to illustrate that our watermark is hard to decrypt. Lastly, our detection process is agnostic and does not rely on any original prompt and LLMs.
@@ 2 Related Work
Generated text watermarking. Generated text watermarking primarily operates on existing text, altering aspects like format (Brassil et al., 1995; Por et al., 2012; Sato et al., 2023; Rizzo et al., 2016), lexical choices (Yoo et al., 2023; Yang et al., 2023; Munyer & Zhong, 2023; Yang et al., 2021), or syntax (Atallah et al., 2001; Topkara et al., 2006; Meral et al., 2009), while preserving the underlying semantics. EASYMARK (Sato et al., 2023), a format-based watermarking, utilizes subtle changes in Unicode characters and spacing to embed watermarks in text. For word-level modifications, Yang et al. (2021), based on context-aware lexical substitution, utilizes BERT to identify semantically suitable synonym candidates for watermarking. Meral et al. (2009) introduces a morphosyntax-based natural language watermarking scheme that uses syntactic tree diagrams to analyze the structure and functional dependencies of text. In addition to the aforementioned methods that primarily rely on specific rules and might lead to unnatural modifications, some neural-based approaches (Zhang et al., 2023; Abdelnabi & Fritz, 2021) involve encoding text with specific message signatures using a pre-trained language model. Abdelnabi & Fritz (2021) propose a method where both sentence embedding and message signatures embedding, extracted by an encoder, are fed into a decoder to generate the watermarked text
LLMs watermarking. Rather than adding a watermark to pre-existing text, an increasing number of studies are focusing on embedding watermarks directly into the text during the LLM generation process (Wang et al., 2023; Fairoze et al., 2023). Some research efforts are dedicated to enhancing the robustness of watermarks (Kirchenbauer et al., 2023a, b; Liu et al., 2023b; Zhao et al., 2023; Ren et al., 2023), aiming to ensure the persistence of watermarking even after exposure to various attacks. Zhao et al. (2023) introduce a provable robust watermark using a fixed vocabulary ‘green/red’ list generated using a shared random key. Ren et al. (2023) utilize the semantics of the generated text as a hash key to separate the vocabulary, which is achieved by discretizing the continuous embedding space. Some works focus on enhancing the security of watermarks (Hu et al., 2023; Wu et al., 2023; Christ et al., 2023; Liu et al., 2023a), aiming to make them difficult to forge. Hu et al. (2023) propose an unbiased watermark using two re-weighting methods. Liu et al. (2023b) proposed a semantic invariant watermark that balances robustness and security. However, it relies on original prompts for detection, which is often impractical to access. The watermarks mentioned above are inserted into the text by manipulating the logits produced by LLMs. Furthermore, some studies Kuditipudi et al. (2023); Christ et al. (2023) introduce watermarks using specially crafted sampling methods. These methods are unbiased and designed to ensure that the watermark has a minimal impact on the original LLM distributions to improve text quality.
@@
Our work focuses on the LLMs watermarking. Previous works in this field have made significant strides in improving the robustness, text quality, and security aspects of watermarking techniques. However, it remains challenging to consider all these aspects simultaneously. Therefore, we intend to propose a holistic watermark approach that achieves strong robustness and high-quality text generation, while also being difficult to forge.
@@
Notations. Given a Language Model (LM), denote $\mathcal{V}$ as the vocabulary of LM, and denote $\mathcal{|\mathcal{V}|}$ as the size of vocabulary. Given a input prompt $R=\{R^{(1)},R^{(2)},...,R^{(N)}\}$, the LM will generate a sequence of tokens $S=\{S^{(1)},S^{(2)},...,S^{(T)}\}$ based on the prompt. We denote $S_{0:t}=\{S^{(1)},...,S^{(t)}\}$ and $S_{0}=\emptyset$, and $r^{(n)}$ and $s^{(t)}$ are concrete tokens in the vocabulary, i.e., $r^{(n)}$, $s^{(t)}\in\mathcal{V}$. As an auto-regressive model, LM will predict the probability distribution $P_{LM}(\cdot|R,S_{0:t})$ over each token in $\mathcal{V}$ at time step $t+1$ based on the prompt and the preceding sequence, and sample the next token based on this probability distribution until the token of termination is sampled.
@@ 3 Adaptive Watermark Generation
In this section, we present our adaptive watermarking technique designed to minimize the impact on text quality, enhance watermark security, and maintain robustness simultaneously. Section 3.1 introduces Adaptive Watermark Token Identification to ensure watermark robustness while minimizing the impact of text quality in generated content. The Semantic-based Logits Scaling Vector Extraction, as discussed in Section 3.2, replaces the widely used ‘green/red’ list to separate the vocabulary, making it more difficult to forge. In Section 3.3, we propose a new distribution perturbation method to inject watermarking, further improving watermarked text quality.
3.1 Adaptive Watermark Token Identification
@@
The Adaptive Watermark Token Identification (see, Figure 1) is designed to minimize the impact of the watermark on the text quality while maintaining its robustness. The watermarking is added in an adaptive manner, i.e., only the distribution of those tokens with high entropy will be perturbed by our watermark method.
Shannon Entropy  (Shannon, 1948) measures the uncertainty of a distribution, which is defined as $H(P_{X})=-\sum_{x\in\mathcal{X}}P_{X}(x){\log}P_{X}(x)$, where $P_{X}$ is a discrete probability distribution defined over the space $\mathcal{X}$, and $0\leq H(P_{X})\leq\log|\mathcal{X}|$. When the entropy is low, the distribution becomes more concentrated on specific tokens. Conversely, when the entropy is high, the distribution becomes more spread out with larger uncertainty. A high entropy distribution in LLMs means that the original distribution offers multiple reasonable choices during sampling. Thus, perturbing the high entropy distribution to insert certain patterns may only subtly impact the sampling process, ensuring that the sampled tokens are still reasonable in the context.
@@
More specifically, given a prompt $R$, the LLM will generate a logits vector $l^{(t+1)}$ over all the tokens in $\mathcal{V}$ at time step $t+1$. The softmax function is then applied to convert the logits into a probability distribution $P_{LM}(\cdot|R,S_{0:t})$, where the probability of the $k$-th token in the $\mathcal{V}$ is $p^{(t+1)}_{k}={\rm exp}(l^{(t+1)}_{k})/\sum_{i=1}^{|\mathcal{V}|}{\rm exp}(l^{(t+1)}_{i}).$ The LLM will sample a token from this probability distribution as the next token using a sampling method, such as Top-$K$ sampling or Top-$p$ sampling. At different time steps, the randomness and uncertainty of the sampling distribution may vary based on the context of the generated text.
@@
To ensure reliable entropy estimation during text generation and detection process without relying on the watermarked LLM, we use an auxiliary Measurement Model (MM) to compute the entropy of the generated text at each time step. The MM can be an auto-regressive model that is significantly smaller in scale compared to the watermarked LLMs, such as GPT-2 (Radford et al., 2019), OPT-350M (Zhang et al., 2022). The entropy measured with MM is expressed as
@@
We set a predefined entropy threshold, denoted as $\alpha$, and avoid disturbing distributions whose entropy is below this threshold. Therefore, our watermark only perturbs those distributions with $H\left(P_{MM}(\cdot|S_{0:t})\right)\geq\alpha$. We refer to the token sampled from the intact distribution as un-watermarked token and the token sampled from the perturbed distribution as watermarked token. As the threshold $\alpha$ increases, two effects become evident. Firstly, there will be more un-watermarked tokens during the generation process. Secondly, the impact of watermarked tokens diminishes, mainly due to the higher inherent uncertainty in their high-entropy distributions.
In addition, due to the fact that only watermarked tokens are involved in the detection process as described in Section 4, adaptive watermarking also ensures robustness by filtering out un-watermarked tokens.

### v2

@@ Abstract
The advancement of Large Language Models (LLMs) has led to increasing concerns about the misuse of AI-generated text, and watermarking LLM-generated text has emerged as a potential solution. However, it is challenging to generate high-quality watermarked text while maintaining robustness, security, and the ability to detect watermarks without prior knowledge of the prompt and model. This paper proposes an adaptive text watermarking strategy to address such a challenge. To improve the text quality and maintain robustness, we adaptively add watermarking to token distributions with high entropy measured by an auxiliary model and keep the low-entropy token distributions untouched. For the sake of security and to further minimize the watermark’s impact on text quality, instead of using a fixed green/red list generated from a random secret key, which can be vulnerable to decryption and forgery, we adaptively scale up the output logits based on the semantic embedding of previously generated text using a well designed semantic mapping model. Our experiments involving various LLMs demonstrate that our approach achieves comparable robustness performance to existing watermark methods. Additionally, the text generated by our method has perplexity comparable to that of un-watermarked LLMs while maintaining sufficient security.111 Our code is available at https://github.com/yepengliu/adaptive-text-watermark.git.
@@ 1 Introduction
The rapid development of Large Language Models (LLMs) (Zhang et al., 2022; Touvron et al., 2023) has ushered in transformative possibilities but also brings forth potential challenges. One primary challenge lies in the misuse of AI-generated text for malicious purposes, such as spreading disinformation, generating fake news, or engaging in plagiarism. The sheer volume and quality of AI-generated text make it increasingly difficult to discern between human and AI-generated content. As a solution, implementing a text watermark on LLMs proves to be an effective method for detecting AI-generated text.
@@
Specifically, the watermarking technique encodes certain patterns into the text generated by LLMs, which are imperceptible to humans but can be detected by the algorithm. This is usually achieved by introducing well-designed perturbations to the original logits or using a special sampling method. Kirchenbauer et al. (2023a) design a watermark for LLM by randomly separating the vocabulary into ‘green’ and ‘red’ lists and increasing the frequency of tokens from the ‘green’ list. The previously generated token is used as a secret key to determine the ‘green/red’ list for the next token. This method achieves good detection performance when the watermarked text remains unmodified. However, after modifications such as word replacement and paraphrasing, this method will fail as the ‘green/red’ list cannot be reliably recovered from the modified text. As it is common for individuals to alter AI-generated text rather than use it directly, improving the robustness of the watermark becomes an essential requirement.
@@
To improve the robustness of LLM watermarking, Zhao et al. (2023) propose to use a fixed ‘green/red’ list for each token, which is generated from a random secret key shared with the detector. Both empirical and theoretical evidence demonstrate that this method substantially enhances robustness. However, more recently, Sadasivan et al. (2023) introduced a spoofing attack specifically designed to deduce hidden watermark patterns by analyzing the token frequency, thereby enabling the forgery of watermarked text. Such an attack offers a method for adversaries to circumvent the watermark, potentially damaging the reputation of the watermarking system. Regrettably, it has been shown that the algorithm proposed in Zhao et al. (2023) is vulnerable to spoofing attack, and one can infer the fixed ‘green’ token list and forge the watermarked text.
@@
The aforementioned papers enhance the frequency of tokens in the ‘green’ list during generation by uniformly adding a fixed value to the logits of those ‘green’ tokens. This way of perturbing the distribution may change the distribution significantly, deteriorating the quality of the watermarked text. To tackle this issue, Kuditipudi et al. (2023) presents a distortion-free watermarking method that maintains the original probability distribution of LLMs and embeds the watermark during the token sampling process. However, such a distortion-free watermarking lacks sufficient robustness against modifications, like paraphrasing.
@@
Inspired by existing research, designing an effective watermark algorithm requires us to consider at least the following three factors comprehensively. (1) Robustness: The watermark should remain detectable after modifications without significantly changing the semantics of the watermarked text. (2) Security: The watermark should be resilient against efforts to reverse-engineer or decryption without authorization. (3) Text quality: LLMs-generated watermarked text should preserve a level of perplexity comparable to the un-watermarked text. Among existing works, achieving high-quality text often comes with the trade-off of robustness, and prioritizing strong robustness can potentially compromise security. Therefore, designing a watermarking scheme that simultaneously guarantees robustness, security, and text quality poses a significant challenge.
@@
Figure 1: Workflow of the proposed Adaptive Watermark. Our Adaptive Watermark method will assess the entropy of the distribution for all previously generated tokens and add watermarks only to high-entropy tokens. This figure illustrates two cases in the text generation process. For the first example, the distribution of the next token has high entropy as measured by AWTI, indicating high uncertainty. Then, the SLSVE module will extract the logits scaling vector based on the semantics of this text. Subsequently, we apply AWTS to perturb the sampling distribution. As for the second text, it results in low entropy measured by AWTI, suggesting a low uncertainty for the next token. Then, the next token will be directly sampled from the original probability distribution.
@@
In this work, our goal is to provide a holistic approach that integrates robustness and security with the production of high-quality text. Additionally, we ensure the watermark can be detected in an agnostic manner, meaning that the watermark detection process is independent of the original LLMs and prompts. Our watermark strategy consists of three components, as shown in Figure 1.
@@
  • To minimize the impact of the watermark on the text quality while maintaining the robustness, we propose the Adaptive Watermark Token Identification (AWTI) so that the watermarking is added adaptively to high-entropy token distributions and leaves the low-entropy token distributions intact. This method employs an auxiliary Measurement Model to assess the entropy of each token, ensuring that the entropy can be reliably estimated without access to the original LLMs and prompts.
@@
  • To further enhance the security of our method, we introduce the Semantic-based Logits Scaling Vector Extraction (SLSVE). Our logits scaling vector is obtained by passing the semantics of the generated text (using a pre-trained Sentence Embedding Model) to a Semantic Mapping Model, which is hard to reverse-engineer.
@@
We conduct extensive experiments over various language models (OPT-6.7B (Zhang et al., 2022), GPT-J-6B (Wang & Komatsuzaki, 2021), and Mistral-7B (Jiang et al., 2023)) to show the effectiveness of our method. We employ perplexity as a metric to evaluate text quality, and our results indicate that the quality of the watermarked text remains comparable to that of the un-watermarked text. To show the robustness of our method, we paraphrase the watermarked text using GPT3.5 and DIPPER (Krishna et al., 2023). The outcomes are on par with those achieved by Zhao et al. (2023), which is the most robust method in the literature. The spoofing attack is used to illustrate that our watermark is hard to decrypt. Lastly, our detection process is agnostic and does not rely on any original prompt and LLMs.
@@ 2 Related Work
Generated text watermarking primarily operates on existing text, altering aspects like format (Brassil et al., 1995; Por et al., 2012; Sato et al., 2023; Rizzo et al., 2016), lexical choices (Yoo et al., 2023; Yang et al., 2023; Munyer & Zhong, 2023; Yang et al., 2022), or syntax (Atallah et al., 2001; Topkara et al., 2006; Meral et al., 2009), while preserving the underlying semantics. EASYMARK (Sato et al., 2023), a format-based watermarking, utilizes subtle changes in Unicode characters and spacing to embed watermarks in text. For word-level modifications, Yang et al. (2022), based on context-aware lexical substitution, utilizes BERT to identify semantically suitable synonym candidates for watermarking. Meral et al. (2009) introduces a morphosyntax-based natural language watermarking scheme that uses syntactic tree diagrams to analyze the structure and functional dependencies of text. In addition to the aforementioned methods that primarily rely on specific rules and might lead to unnatural modifications, some neural network-based approaches (Zhang et al., 2023b; Abdelnabi & Fritz, 2021) add watermarking by regenerating the text with specific message signatures using a pre-trained language model, while keeping the semantics unchanged.
LLMs watermarking. Rather than adding a watermark to existing text, an increasing number of studies are focusing on embedding watermarks directly into the text during the LLM generation (Wang et al., 2023; Fairoze et al., 2023; Hu et al., 2023; Lee et al., 2023; Huo et al., 2024; Fu et al., 2024; Zhao et al., 2024; Zhang et al., 2023a; Tu et al., 2023).
Some research efforts are dedicated to enhancing the robustness of watermarks (Kirchenbauer et al., 2023a, b; Liu et al., 2023b; Zhao et al., 2023; Ren et al., 2023), aiming to ensure the persistence of watermarking even after exposure to various attacks. Zhao et al. (2023) introduce a provable robust watermark using a fixed vocabulary ‘green/red’ list generated using a shared random key. Ren et al. (2023) utilize the semantics of the generated text as a hash key to separate the vocabulary, which is achieved by discretizing the continuous embedding space. Hou et al. (2023) propose a sentence-level semantic watermarking technique that utilizes the semantics of previously generated sentences to divide the semantic space into “valid” and “blocked” regions and generates new sentences through rejection sampling until its semantics reside within the “valid” region.
Some works focus on enhancing the security of watermarks (Wu et al., 2023; Christ et al., 2023; Liu et al., 2023a), aiming to make them difficult to forge. Liu et al. (2023b) proposed a semantic invariant watermark that balances robustness and security. However, this method relies on the original prompts for detection, which are often impractical to access. Additionally, the way it perturbs the distribution results in noticeably higher text perplexity compared to un-watermarked text.
All the watermarks mentioned above are inserted into the text by manipulating the logits produced by LLMs, which inevitably introduces some distortion in text generation. Some studies Kuditipudi et al. (2023); Christ et al. (2023) introduce watermarks using specially crafted sampling methods. These methods are designed to ensure that the watermark has a minimal distortion on the original LLM distributions to improve text quality.
@@
Our work focuses on the LLMs watermarking. Previous works in this field have significantly improved the robustness, text quality, and security aspects of watermarks in a separate manner. However, it remains challenging to consider all these aspects simultaneously. Therefore, we intend to propose a holistic watermark approach that achieves strong robustness and high-quality text generation while also being difficult to forge.
@@
Notations. Given a Language Model (LM), denote $\mathcal{V}$ as the vocabulary of LM, and denote $\mathcal{|\mathcal{V}|}$ as the size of vocabulary. For a input prompt $R=\{R^{(1)},R^{(2)},...,R^{(N)}\}$, the LM will generate a sequence of tokens $S=\{S^{(1)},S^{(2)},...,S^{(T)}\}$ based on the prompt. We denote $S_{0:t}=\{S^{(1)},...,S^{(t)}\}$ and $S_{0}=\emptyset$, and $r^{(n)}$ and $s^{(t)}$ are concrete tokens in the vocabulary, i.e., $r^{(n)}$, $s^{(t)}\in\mathcal{V}$. As an auto-regressive model, LM will predict the probability distribution $P_{LM}(\cdot|R,S_{0:t})$ over each token in $\mathcal{V}$ at time step $t+1$ based on the prompt and the preceding sequence, and sample the next token based on this probability distribution until the token of termination is sampled.
@@ 3 Adaptive Watermark Generation
In this section, we present our adaptive watermarking technique designed to minimize the impact on text quality, enhance security, and maintain robustness simultaneously. Section 3.1 introduces Adaptive Watermark Token Identification to ensure watermark robustness while minimizing the impact of text quality in generated content. The Semantic-based Logits Scaling Vector Extraction, as discussed in Section 3.2, generates different logits scaling vectors based on the semantics of the generated text, which adaptively perturb the distribution of the LLM. In Section 3.3, we propose a new distribution perturbation method to inject watermarking, further improving watermarked text quality.
3.1 Adaptive Watermark Token Identification (AWTI)
@@
The Adaptive Watermark Token Identification (see Figure 1-➀) is designed to minimize the impact of the watermark on the text quality while maintaining its robustness by reducing the percentage of perturbed tokens during the text generation process. The watermarking is added in an adaptive manner, i.e., only the distribution of those tokens flagged by our auxiliary language model as having high entropy will be perturbed by our watermarking method.
@@
Specifically, given a prompt $R$, the LLM will generate a logits vector $l^{(t+1)}$ over all the tokens in $\mathcal{V}$ at time step $t+1$. The softmax function is then applied to convert the logits into a probability distribution $P_{LM}(\cdot|R,S_{0:t})$, where the probability of the $k$-th token in the $\mathcal{V}$ is $p^{(t+1)}_{k}={\rm exp}(l^{(t+1)}_{k})/\sum_{i=1}^{|\mathcal{V}|}{\rm exp}(l^{(t+1)}_{i}).$ The LLM will sample a token from this probability distribution using methods such as Top-$K$ sampling or Top-$p$ sampling. As shown in the example of Figure 1, the sampling distribution may vary based on the context of the generated text at different steps.
Shannon Entropy  (Shannon, 1948) measures the uncertainty of a distribution, which is defined as $H(P_{X})\triangleq-\sum_{x\in\mathcal{X}}P_{X}(x){\log}P_{X}(x)$, where $P_{X}$ is a discrete probability distribution defined over the space $\mathcal{X}$, and $0\leq H(P_{X})\leq\log|\mathcal{X}|$. When the entropy is low, the distribution becomes more concentrated on specific tokens. Conversely, when the entropy is high, the distribution becomes more spread, indicating larger uncertainty. A high entropy distribution in LLMs means that the original distribution offers multiple reasonable choices during text generation. Thus, inserting certain patterns by only perturbing the high entropy distribution will minimize the impact of the sampling process, ensuring that the generated tokens remain contextually appropriate.
We note that Lee et al. (2023) also discusses the strategy of perturbing only the token with high entropy distributions for code generation. Their approach, however, requires access to both the original LLM and prompts, rendering it impractical for watermarking detection. Additionally, utilizing the original LLM for detection is not only time-consuming but also incurs substantial computational costs.
@@
To ensure reliable entropy estimation during the text generation and detection process without relying on the watermarked LLM or the original prompts, we use an auxiliary Measurement Model (MM) to compute the entropy of the generated text at each time step. The entropy measured with MM is expressed as
@@
The MM can be a language model that is significantly smaller in scale compared to the watermarked LLMs, such as GPT-2 (Radford et al., 2019) or OPT-350M (Zhang et al., 2022). We use a predefined entropy threshold, denoted as $\alpha$, and our watermark only perturbs those distributions with $H\left(P_{MM}(\cdot|S_{0:t})\right)\geq\alpha$. We refer to the token sampled from the intact distribution as un-watermarked token and the token sampled from the perturbed distribution as watermarked token. As the threshold $\alpha$ increases, two effects become evident. Firstly, there will be more un-watermarked tokens during the generation process. Secondly, the impact of each watermarked token diminishes, mainly due to the higher inherent uncertainty in their distributions. As a result, both effects will enhance the quality of the watermarked text.
It is important to acknowledge that the entropy computed by the auxiliary MM may differ from that of the original watermarked LLM, but this discrepancy does not affect the efficacy of our watermarking. The auxiliary MM is employed primarily to assist in identifying potential watermarked tokens during the detection phase, eliminating the need for the original watermarked LLM and prompt. As long as we can identify watermarked tokens in the detection phase using such entropy estimates, the performance will not be affected. In addition, due to the fact that only watermarked tokens are involved in the detection process as described in Section 4, adaptive watermarking may also benefit robustness by filtering out un-watermarked tokens.